Jupiter's average distance to the Sun is 7.784×1011m. How much time does it take for Jupiter to complete one orbit around the Sun?
Answer #1

here,

the average distance from SUN to Jupiter , r = 7.784 * 10^11 m

the mass of SUN , m = 2 * 10^30 kg

the time taken , T = 2*pi*sqrt(r^3 /(G * m))

T = 2*pi * sqrt((7.784 * 10^11)^3 /(6.67 * 10^-11 * 2 * 10^30))

T = 3.73 * 10^7 s

the time period of orbit is 3.73 * 10^7 s